# A convex formulation of covariate-adjusted Gaussian graphical models via natural parametrization

### v1

@@ 4 Theoretical Properties
In this section we will analyze the estimation error and support recovery of (15). For two sequences of real numbers $a_{n}$ and $b_{n}$, we write $a_{n}\precsim b_{n}$ if $a_{n}=O(b_{n})$, i.e. there exists constants $C>0$ and $N>0$ so that $a_{n}<Cb_{n}$ for all $n\geq N$. If $a_{n}\precsim b_{n}$ and $b_{n}\precsim a_{n}$, we write $a_{n}\asymp b_{n}$. We write $a_{n}=o(b_{n})$ if $\lim_{n\rightarrow\infty}a_{n}/b_{n}=0$. With the understanding that $j$ is fixed, we will suppress the subscript $j$ when referring to $\boldsymbol{\mathbf{\gamma}}_{j}$, $\boldsymbol{\mathbf{\beta}}_{j}$ and $\boldsymbol{\mathbf{\varepsilon}}_{j}$ for notational convenience.
Let $(\hat{\boldsymbol{\mathbf{\gamma}}},\hat{\boldsymbol{\mathbf{\beta}}})$ be the solution to (15) and let $(\boldsymbol{\mathbf{\gamma}},\boldsymbol{\mathbf{\beta}})$ be the true parameters. Denote by $S_{\gamma}$ and $S_{\beta}$ the support sets of $\boldsymbol{\mathbf{\gamma}}$ and $\boldsymbol{\mathbf{\beta}}$ respectively. Further let $S_{\beta,g}$ index the active groups of $\boldsymbol{\mathbf{\beta}}$, that is $S_{\beta,g}=\left\{h:\boldsymbol{\mathbf{\beta}}_{j,h}\neq\boldsymbol{\mathbf{0}},\,h\in[q]\right\}$. Denote by $s_{\gamma},\,s_{\beta}$ and $s_{\beta,g}$ the cardinalities of these sets. For a square matrix $\boldsymbol{\mathbf{M}}$, denote by $\lambda_{\mathrm{min}}(\boldsymbol{\mathbf{M}})$ and $\lambda_{\mathrm{max}}(\boldsymbol{\mathbf{M}})$ its minimum and maximum eigenvalue, respectively.
We require the following assumptions on the model (1) with specification (7).
@@
for some constants $0<\phi_{0}\leq\phi_{1}<\infty$. Furthermore, there exists a constant $M>0$ such that $\lvert\boldsymbol{\mathbf{u}}^{(i)}_{j}\rvert\leq M$ for all $j\in[q]$ and $i\in[n]$.
Assumption 1 is the same as Assumptions 1 and 5 in Zhang and Li [2022] without requiring a bound on $\lVert\boldsymbol{\mathbf{\beta}}\rVert_{1}$. The boundedness of the covariates is not restrictive in the eQTL setting since SNPs are often binary coded.
Denote by $[\boldsymbol{\mathbf{U}},\boldsymbol{\mathbf{W}}_{-j}]$ the $n\times(p(q+1)-1)$ matrix that is the concatenation of the covariate matrix $\boldsymbol{\mathbf{U}}$ with the interaction matrix $\boldsymbol{\mathbf{W}}_{-j}$. Define the following Gram matrix of covariates, responses, and interactions:
$\boldsymbol{\mathbf{\Sigma}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}=\operatorname{\operatorname{\mathbb{E}}}\left(\frac{[\boldsymbol{\mathbf{U}},\boldsymbol{\mathbf{W}}_{-j}]^{\top}[\boldsymbol{\mathbf{U}},\boldsymbol{\mathbf{W}}_{-j}]}{n}\right).$
Our next assumption bounds the eigenvalues of $\boldsymbol{\mathbf{\Sigma}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}$, which is needed to characterize the joint distribution of $\boldsymbol{\mathbf{u}}^{(i)}$, $\boldsymbol{\mathbf{x}}^{(i)}$, and their interactions.
@@
We assume there exist positive constants $m_{0},M_{0}$ such that
$m_{0}\leq\lambda_{\mathrm{min}}(\boldsymbol{\mathbf{\Sigma}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}})\leq\lambda_{\mathrm{max}}(\boldsymbol{\mathbf{\Sigma}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}})\leq M_{0}.$
Elementwise boundedness in Assumption 1 implies that $\boldsymbol{\mathbf{u}}^{(i)}$ is elementwise sub-Gaussian. Our theoretical analysis further requires that our design matrix $[\boldsymbol{\mathbf{U}},\boldsymbol{\mathbf{W}}_{-j}]$ is elementwise sub-Gaussian. Assumption 3 ensures this since each entry of $\boldsymbol{\mathbf{W}}_{-j}$ is the product of a sub-Gaussian and a bounded random variable.
@@
The marginal distribution of $\bm{X}$ is elementwise sub-Gaussian with bounded sub-Gaussian norm.
The following assumption controls how the sparsity of the true parameters may scale with the sample size $n$.
@@
Let $p$ and $q$ be the number of responses and covariates, respectively. The true sparsities $s_{\gamma}$ and $s_{\beta}$ satisfy
  1. 1.
    $(s_{\gamma}+s_{\beta})\sqrt{\log(pq)}=o(\sqrt{n})$,
  2. 2.
    $(s_{\gamma}+s_{\beta})\log(pq)=o(n/\log n)$.
In practice we consider estimates $(\hat{\boldsymbol{\mathbf{\gamma}}},\hat{\boldsymbol{\mathbf{\beta}}})$ of varying sparsities and select from this pool of candidate models via cross-validation. Define $\hat{s}_{\gamma}^{\mathrm{max}}$ and $\hat{s}_{\beta}^{\mathrm{max}}$ to be the maximum sparsity of $\boldsymbol{\mathbf{\gamma}}$ and $\boldsymbol{\mathbf{\beta}}$ out of all candidate models, chosen so that $s_{\gamma}<s_{\gamma}^{\mathrm{max}}$ and $s_{\beta}<s_{\beta}^{\mathrm{max}}$. Our first theorem describes the $\ell_{2}$ estimation error of the nodewise solution.
@@
$\lambda=C\frac{\sigma_{\varepsilon}}{\sqrt{n}}\left(\frac{s_{\gamma}\log(eq/s_{\gamma})+s_{\beta}\log(ep)}{s_{\gamma}+s_{\beta}}+\frac{s_{\beta,g}}{s_{\gamma}+s_{\beta}}\log(eq/s_{\beta,g})\right)^{1/2}\quad\text{and}\quad\lambda_{g}=\lambda\sqrt{\frac{s_{\gamma}+s_{\beta}}{s_{\beta,g}}}$ (17)
we have
$\lVert\hat{\boldsymbol{\mathbf{\gamma}}}-\boldsymbol{\mathbf{\gamma}}\rVert_{2}^{2}+\lVert\hat{\boldsymbol{\mathbf{\beta}}}-\boldsymbol{\mathbf{\beta}}\rVert_{2}^{2}\precsim\frac{\sigma_{\varepsilon}^{2}}{n}\left(s_{\gamma}\log(eq/s_{\gamma})+s_{\beta}\log(ep)+s_{\beta,g}\log(eq/s_{\beta,g})\right).$
with probability at least $1-C_{1}\exp\left\{-C_{2}(s_{\gamma}\log(eq/s_{\gamma})+s_{\beta}\log(ep)+s_{\beta,g}\log(eq/s_{\beta,g}))\right\}$ where $C$, $C_{1}$, and $C_{2}$ are positive constants.
The bound on the sparsity of candidate models helps to bound the errors in Theorem 1 and is also assumed in Zhang and Li [2022].
Our second result proves that our method achieves support recovery with high probability under some additional conditions. Denote the $(k,\ell)$-th entry of $\boldsymbol{\mathbf{\Sigma}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}$ by $\boldsymbol{\mathbf{\Sigma}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}(k,\ell)$.
@@
Define $\tau_{j}=1+\sqrt{(s_{\gamma}+s_{\beta})/s_{\beta,g}}$. We assume that there exists a constant $c_{0}>2/m_{0}$ such that
$\max_{k\neq\ell}\lvert\boldsymbol{\mathbf{\Sigma}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}(k,\ell)\rvert\leq\frac{1}{c_{0}(1+8\tau_{j})(s_{\beta}+s_{\gamma})}.$
@@
Suppose Assumptions 1-5 hold. Additionally, suppose $\log p\asymp\log q$ and that
@@
for some constant $A_{1}>0$. Then with $\lambda$ and $\lambda_{g}$ defined as in Theorem 1, we have
$\max\left\{\lVert\hat{\boldsymbol{\mathbf{\gamma}}}-\boldsymbol{\mathbf{\gamma}}\rVert_{\infty},\lVert\hat{\boldsymbol{\mathbf{\beta}}}-\boldsymbol{\mathbf{\beta}}\rVert_{\infty}\right\}\leq\frac{3}{m_{0}}(\lambda+\lambda_{g})\left(1+\frac{6(1+4\tau_{j})^{2}}{(1+8\tau_{j})(c_{0}m_{0}-2)}\right)$
with probability at least $1-C_{3}\exp(-C_{4}\log p)$ for some positive constants $C_{3},C_{4}$.
Theorem 2 implies that our method achieves support recovery with high probability if $\boldsymbol{\mathbf{\gamma}}$ and $\boldsymbol{\mathbf{\beta}}$ satisfy a minimal signal strength condition.
@@
$\kappa_{0}$ $=\frac{3}{m_{0}}(\lambda+\lambda_{g})\left(1+\frac{6(1+4\tau_{j})^{2}}{(1+8\tau_{j})(c_{0}m_{0}-2)}\right),$
$\tilde{S}_{\gamma}$ =$=\left\{k:\lvert\hat{\gamma}_{k}\rvert>\kappa_{0}\right\},\quad\text{and}\quad\tilde{S}_{\beta}=\left\{k:\lvert\hat{\beta}_{k}\rvert>\kappa_{0}\right\}.$
If it holds that
@@
we have that $\operatorname{\operatorname{\mathbb{P}}}(S_{\gamma}=\tilde{S}_{\gamma}\;\text{and}\;S_{\beta}=\tilde{S}_{\beta})\geq 1-C_{3}\exp(-C_{4}\log p)$.
Zhang and Li [2022] provides two theorems related to the estimation error of $\hat{\boldsymbol{\mathbf{\beta}}}$; an oracle rate where the true $\boldsymbol{\mathbf{\Gamma}}$ in (3) is known and a rate where $\hat{\boldsymbol{\mathbf{\Gamma}}}$ is estimated via independent regressions in stage 1 of the procedure. It is shown that the oracle rate, given by
@@
$\lVert\hat{\boldsymbol{\mathbf{\beta}}}-\boldsymbol{\mathbf{\beta}}\rVert_{2}^{2}\precsim\frac{\sigma^{2}_{\varepsilon_{j}}}{n}(s_{\beta}\log(ep)+s_{\beta,g}\log(eq/s_{\beta,g})),$
can be achieved by the two-stage rate under the assumptions $\log(pq)=O(n^{1/6})$, $s_{\beta}=o(n^{1/6})$, and $s_{\gamma}=o(n^{1/3})$. By comparison, Theorem 1 shows that our method achieves the oracle rate under milder assumptions; more explicitly, Assumption 4 is satisfied when $\log(pq)=O(n^{1/3})$ and $s_{\gamma}+s_{\beta}=o(n^{1/3})$. The milder scaling assumptions in $\log(pq)$ and $s_{\beta}$ in our formulation stems from the fact that (15) is jointly convex in $(\boldsymbol{\mathbf{\gamma}},\boldsymbol{\mathbf{\beta}})$, allowing for their simultaneous estimation. By comparison, a more stringent scaling assumption is needed to control the estimation error of $\boldsymbol{\mathbf{\gamma}}_{j}$ incurred in stage 1 of Zhang and Li [2022]. These comments apply as well to the support recovery results in Theorem 2.
5 Simulation Study
In this section, we compare our method Covariate-adjusted Sparse Precision with Natural Estimation (cspine) with the Gaussian graphical model regression method developed in Zhang and Li [2022] (RegGMM) under extensive simulations. We are interested in the behavior of the two methods under data generated according to both models.

### v2

@@ 4 Theoretical properties
In this section, we present theoretical guarantees for the estimator defined in (14), which characterize the accuracy of the recovered graph structure and its dependence on covariates. We begin by introducing notation and stating assumptions on the data-generating process. We then present our main results, including $\ell_{2}$ estimation error bounds and support recovery guarantees, followed by a discussion comparing our results with those of zhangli.
For two sequences of real numbers $a_{n}$ and $b_{n}$, we write $a_{n}\precsim b_{n}$ if $a_{n}=O(b_{n})$, i.e. there exists constants $C>0$ and $N>0$ so that $a_{n}<Cb_{n}$ for all $n\geq N$. If $a_{n}\precsim b_{n}$ and $b_{n}\precsim a_{n}$, we write $a_{n}\asymp b_{n}$. We write $a_{n}=o(b_{n})$ if $\lim_{n\rightarrow\infty}a_{n}/b_{n}=0$. Our analysis holds for any $j\in[p]$. For notational simplicity, we suppress the subscript $j$ when referring to $\boldsymbol{\mathbf{\gamma}}_{j}$, $\boldsymbol{\mathbf{\beta}}_{j}$, $\boldsymbol{\mathbf{\varepsilon}}_{j}$ and $\boldsymbol{\mathbf{W}}_{-j}$ in this section.
[…]
@@
Suppose $\{\boldsymbol{\mathbf{u}}^{(i)}\}_{i=1}^{n}$ are i.i.d. mean zero random vectors with uniformly bounded entries: there exists a constant $M_{u}>0$ such that $\lvert u^{(i)}_{h}\rvert\leq M_{u}$ for all $h\in[q]$ and $i\in[n]$.
@@
The responses $\boldsymbol{\mathbf{x}}^{(i)}$ are elementwise sub-Gaussian with bounded sub-Gaussian norm: there exists a constant $K_{X}>0$ such that $\lVert x^{(i)}_{j}\rVert_{\psi_{2}}\leq K_{X}$ for all $j\in[p]$ and $i\in[n]$.
Let $\boldsymbol{\mathbf{U}},\boldsymbol{\mathbf{W}}]\in\mathbb{R}^{n\times(p(q+1)-1)}$ denote the augmented design matrix in (14). Define the population Gram matrix $\boldsymbol{\mathbf{\Sigma}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}=\operatorname{\operatorname{\mathbb{E}}}\left([\boldsymbol{\mathbf{U}},\boldsymbol{\mathbf{W}}]^{\top}[\boldsymbol{\mathbf{U}},\boldsymbol{\mathbf{W}}]/n\right)$.
@@
There exist constants $0<\phi_{0}^{\ast}\leq\phi_{1}^{\ast}<\infty$ such that
$\phi_{0}^{\ast}\leq\lambda_{\mathrm{min}}(\boldsymbol{\mathbf{\Sigma}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}})\leq\lambda_{\mathrm{max}}(\boldsymbol{\mathbf{\Sigma}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}})\leq\phi_{1}^{\ast}.$
@@
The sparsity levels satisfy
$(s_{\gamma}+s_{\beta})\log(pq)=o(\sqrt{n}/\log n).$
Assumption 12 controls the tail behavior of the covariates, responses, and their pairwise interactions, thus facilitating their concentration behavior. Similar assumptions are also imposed in zhangli. Assumption 3 ensures the augmented design matrix is well-conditioned, preventing excessive collinearity among predictors in the nodewise regression (14). Assumption 3 is closely related to the bounded eigenvalue assumptions on $\operatorname{Cov}(\bm{U})$ and $\operatorname{Cov}(\bm{X}\>|\>\bm{U})$ in zhangli. In Section S5, we provide interpretable sufficient conditions under which Assumption 3 holds. Finally, Assumption 4 governs the scaling of sparsity levels with the sample size. Combined with Theorem 4.1 of kuchi, it allows us to control sparse deviations between sample and population Gram matrices (see Lemmas 5 and 7 in Appendix S4). A sufficient condition for Assumption 4 is $\log(pq)=O(n^{1/6})$ and $s_{\gamma}+s_{\beta}=o(n^{1/6})$, as assumed in zhangli.
[…]
@@
$\lambda=C\frac{\sigma_{\varepsilon}}{\sqrt{n}}\left(\frac{s_{\gamma}\log(eq/s_{\gamma})+s_{\beta}\log(ep)}{s_{\gamma}+s_{\beta}}+\frac{s_{\beta,g}}{s_{\gamma}+s_{\beta}}\log(eq/s_{\beta,g})\right)^{1/2},\quad\lambda_{g}=\lambda\sqrt{\frac{s_{\gamma}+s_{\beta}}{s_{\beta,g}}},$ (18)
we have that
$\lVert\hat{\boldsymbol{\mathbf{\gamma}}}-\boldsymbol{\mathbf{\gamma}}\rVert_{2}^{2}+\lVert\hat{\boldsymbol{\mathbf{\beta}}}-\boldsymbol{\mathbf{\beta}}\rVert_{2}^{2}\precsim\frac{\sigma_{\varepsilon}^{2}}{n}\left\{s_{\gamma}\log(eq/s_{\gamma})+s_{\beta}\log(ep)+s_{\beta,g}\log(eq/s_{\beta,g})\right\}$
holds with probability at least $1-C_{1}\exp\left[-C_{2}\{s_{\gamma}\log(eq/s_{\gamma})+s_{\beta}\log(ep)+s_{\beta,g}\log(eq/s_{\beta,g})\}\right]$ where $C$, $C_{1}$, and $C_{2}$ are positive constants.
We next establish support recovery guarantees, under an additional mutual incoherence condition. Denote the $(k,\ell)$-th entry of $\boldsymbol{\mathbf{\Sigma}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}$ by $\boldsymbol{\mathbf{\Sigma}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}(k,\ell)$.
@@
Define $\tau_{j}=1+\sqrt{(s_{\gamma}+s_{\beta})/s_{\beta,g}}$. Suppose
$\max_{k\neq\ell}\lvert\boldsymbol{\mathbf{\Sigma}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}(k,\ell)\rvert\leq\frac{1}{c_{0}(1+8\tau_{j})(s_{\beta}+s_{\gamma})}$
for a constant $c_{0}>2/\phi^{\ast}_{0}$.
Assumption 5 imposes a mutual incoherence condition [Van2009On] on the augmented design matrix. It controls correlations between active and inactive coordinates and is essential for deriving $\ell_{\infty}$ error bounds necessary for support recovery.
@@
Suppose the assumptions of Theorem 1, together with Assumption 5, hold, and that $\log p\asymp\log q$. Further assume that
@@
for some constant $A_{1}>0$. With the same tuning parameters $\lambda$ and $\lambda_{g}$ as in (18), we have that
$\max\left\{\lVert\hat{\boldsymbol{\mathbf{\gamma}}}-\boldsymbol{\mathbf{\gamma}}\rVert_{\infty},\lVert\hat{\boldsymbol{\mathbf{\beta}}}-\boldsymbol{\mathbf{\beta}}\rVert_{\infty}\right\}\leq\frac{3}{\phi^{\ast}_{0}}(\lambda+\lambda_{g})\left(1+\frac{6(1+4\tau_{j})^{2}}{(1+8\tau_{j})(c_{0}\phi^{\ast}_{0}-2)}\right)$
holds with probability at least $1-C_{3}\exp(-C_{4}\log p)$ for some positive constants $C_{3},C_{4}$.
As in the standard lasso regression, with an additional minimal signal strength condition, the $\ell_{\infty}$ bound in Theorem 2 implies that the support of $\boldsymbol{\mathbf{\gamma}}$ and $\boldsymbol{\mathbf{\beta}}$ can be recovered with high probability by thresholding the estimated coefficients at an appropriate level.
@@
$\kappa_{0}=\frac{3}{\phi^{\ast}_{0}}(\lambda+\lambda_{g})\left(1+\frac{6(1+4\tau_{j})^{2}}{(1+8\tau_{j})(c_{0}\phi^{\ast}_{0}-2)}\right).$
Let $\tilde{S}_{\gamma}=\left\{k:\lvert\hat{\gamma}_{k}\rvert>\kappa_{0}\right\}$ and $\tilde{S}_{\beta}=\left\{k:\lvert\hat{\beta}_{k}\rvert>\kappa_{0}\right\}$. If
@@
then $\operatorname{\operatorname{\mathbb{P}}}(S_{\gamma}=\tilde{S}_{\gamma}\;\text{and}\;S_{\beta}=\tilde{S}_{\beta})\geq 1-C_{3}\exp(-C_{4}\log p)$.
4.1 Understanding the theoretical guarantees
We conclude this section with a discussion of the $\ell_{2}$ estimation error and a comparison with the results of zhangli. The same insights apply to the support recovery guarantees in Theorem 2.
In zhangli, two rates of estimation error are established: an oracle rate and a two-stage rate. The oracle rate is achieved when the true $\boldsymbol{\mathbf{\Gamma}}$ in (3) is known, so that the centered responses can be constructed exactly and the estimation in (7) can be carried out directly. The resulting rate is
@@ 4.1 Understanding the theoretical guarantees
$\lVert\hat{\boldsymbol{\mathbf{\eta}}}_{j}-\boldsymbol{\mathbf{\eta}}_{j}\rVert_{2}^{2}\precsim\frac{\sigma^{2}_{\varepsilon_{j}}}{n}(s_{\beta}\log(ep)+s_{\beta,g}\log(eq/s_{\beta,g})),$ (19)
which holds with probability at least $1-C_{1}\exp[-C_{2}\{s_{\beta}\log(ep)+s_{\beta,g}\log(eq/s_{\beta,g})\}]$ for some positive constants $C_{1},C_{2}$.
In practice, however, $\boldsymbol{\mathbf{\Gamma}}$ is unknown and must be estimated. As reviewed in Section 2.2, zhangli employ a two-stage procedure in which $\boldsymbol{\mathbf{\Gamma}}$ is estimated via the misspecified lasso regression
$\hat{\boldsymbol{\mathbf{\gamma}}}_{j}=\operatorname*{arg\,min}_{\boldsymbol{\mathbf{\gamma}}_{j}}\frac{1}{2n}\left\lVert\boldsymbol{\mathbf{x}}_{j}-\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{\gamma}}_{j}\right\rVert_{2}^{2}+\lambda_{1}\lVert\boldsymbol{\mathbf{\gamma}}_{j}\rVert_{1},$ (20)
where $\lambda_{1}\geq 0$, and the resulting estimate is used to construct centered responses $\hat{\boldsymbol{\mathbf{z}}}_{j}=\boldsymbol{\mathbf{x}}_{j}-\boldsymbol{\mathbf{U}}\hat{\boldsymbol{\mathbf{\gamma}}}_{j}$, which are then used in the second-stage regression (7).
Although the resulting two-stage estimator can attain the oracle rate (19), the associated probability bound depends critically on the tuning parameter $\lambda_{1}$ in (20). Specifically, supposing that $\lambda_{1}\propto\sqrt{\kappa_{1}\log q/n}$ for some $\kappa_{1}>0$. Standard lasso results imply that the first-stage estimation error satisfies $\lVert\hat{\boldsymbol{\mathbf{z}}}_{j}-\boldsymbol{\mathbf{z}}_{j}\rVert_{2}^{2}/n\precsim s_{\gamma}\cdot\kappa_{1}\log q/n$ with probability at least $1-3\exp(-\kappa_{1}\log q)$. To ensure that this error does not dominate the second-stage estimation, a union bound must be applied over all $p\times q$ columns of the design matrix. This leads to the overall success probability of at least $1-C_{3}\exp[C_{4}\{\log p-(\kappa_{1}-1)\log q\}]$ for positive constants $C_{3},C_{4}$.
This dependence introduces a nontrivial tradeoff: if $\kappa_{1}$ is too small, the concentration probability deteriorates. If $\kappa_{1}$ is too large, the first-stage penalty becomes overly aggressive, increasing the bias in the first-stage estimator and potentially inflating the overall error. In addition, when $\log p$ dominates $\log q$, the success probability may no longer decay to zero, further weakening the theoretical guarantees.
By contrast, our estimator matches the oracle rate without requiring a separate centering stage, thereby avoiding error propagation from the first stage and eliminating the need to tune an additional penalty parameter. Furthermore, the oracle rate in zhangli is established under the scaling conditions $\log(pq)=O(n^{1/6})$ and $s_{\gamma}+s_{\beta}=o(n^{1/6})$. In comparison, Assumption 4 is strictly weaker and allows for a wider range of scaling regimes. It permits $\log(pq)$ and $s_{\gamma}+s_{\beta}$ to grow at a faster rates relative to $n$, e.g., $\log(pq)=O(n^{1/4}/\log(n))$ and $s_{\gamma}+s_{\beta}=o(n^{1/4})$, while still ensuring consistency. Note however that the scaling in zhangli may be tightened beyond what is reported.
These theoretical advantages in our framework stem from the joint convexity of (14), which enables simultaneous estimation of $(\boldsymbol{\mathbf{\gamma}},\boldsymbol{\mathbf{\beta}})$ without intermediate approximation steps. In contrast, the two-stage procedure in zhangli requires additional control over the first-stage estimation error, leading to more restrictive conditions. We provide additional empirical evidence of these advantage in Section 5, particularly in settings where accurate first stage estimation in zhangli is challenging.
5 Simulation studies
We compare our method, termed cspine, with zhangli (referred to as RegGMM) through extensive simulation studies. Comparisons with other methods considered in zhangli are omitted, as RegGMM is shown therein to outperform these alternatives. We use the R implementation of RegGMM, which is generously shared by the authors. Our R implementation uses the sparsegl package to solve (14) [liang_sparsegl_2024].
5.1 Data generation